OpenStudy (anonymous):

2x/5 +2=8 showingall work

OpenStudy (anonymous):

subtract the 2: 2x/5=6 multiply by 5: 2x=30 solve

OpenStudy (anonymous):

2x/5 +2=8 2x+10=40 2x=30 x=15